# JEL classification: C11, C32, C55, E50

### v1

@@ Abstract
We propose a high-dimensional structural vector autoregression framework capable of accommodating a large number of linear inequality restrictions on impact impulse responses, structural shocks, and their element-wise products. Combining impact- and shock-inequality restrictions can be flexibly used to sharpen inference and to disentangle structurally interpretable shocks through sign and/or shock constraints. To estimate the model, we develop a highly efficient sampling algorithm that scales well with model dimension and the number of inequality restrictions on impact responses, as well as structural shocks. It remains computationally feasible even when existing algorithms may break down. To demonstrate the practical utility of our approach, we identify five structural shocks and examine the dynamic responses of thirty macroeconomic variables, highlighting the model’s flexibility and feasibility in complex empirical settings. We provide empirical evidence that financial shocks are the most important driver of the dynamics of the business cycle.
@@ 1 Introduction
Uhlig, (2005) popularizes the use of sign restrictions on impulse response functions, offering a flexible and theory-consistent approach to identifying structural shocks in Vector Autoregression (VAR) models. Instead of relying on often implausible zero restrictions, sign restrictions constrain the signs of impulse responses based on economic intuition. Building on this framework, Antolín-Díaz and Rubio-Ramírez, (2018) introduce narrative restrictions, which incorporate external prior information about specific historical episodes-such as well-documented policy interventions or economic events-to further sharpen set identification. By conditioning structural shocks on particular time periods, narrative restrictions enable researchers to embed economically meaningful prior knowledge directly into the model, enhancing interpretability and credibility of the identified shocks. Using narrative restrictions have become popular recently. [^1] Important examples that use narrative restrictions include Zeev, (2018), Altavilla et al., (2019), Furlanetto and Robstad, (2019), Cheng and Yang, (2020), Kilian and Zhou, (2020), Kilian and Zhou, (2022), Laumer, (2020), Redl, (2020), Zhou, (2020), Antolin-Diaz et al., (2021), Caggiano et al., (2021), Larsen, (2021), Ludvigson et al., (2021), Maffei-Faccioli and Vella, (2021), Berger et al., (2022), Fanelli and Marsi, (2022), Inoue and Kilian, (2022), Badinger and Schiman, 2023a , Berthold, (2023), Caggiano and Castelnuovo, (2023), Conti et al., (2023), Harrison et al., (2023), Herwartz and Wang, (2023) Neri, (2023), Reichlin et al., (2023), Ascari et al., (2023), Boer et al., (2024) and Rüth and Van der Veken, (2023). However, in most applications researcher use rather small dimensional VARs with narrative restrictions. Furthermore, the implementation of narrative restriction rely on accept-and-reject algorithms (see, e.g., Antolín-Díaz and Rubio-Ramírez, (2018) or Giacomini et al., (2021)), which may become computationally burdensome as the number of restrictions increases. Therefore, applications with narrative restriction are limited to VARs with a small number of narrative restrictions, typically for just a single structural shock.
@@
In this paper, we propose a high-dimensional Structural Vector Autoregression (SVAR) framework capable of accommodating a large number of linear inequality constraints on impact impulse responses, structural shocks, and their element-wise products. While previous literature has primarily focused on small narrative SVAR models, our approach extends these methods by enabling the inclusion of a much larger set of restrictions as well as a greater number of variables. The ability to incorporate a broader set of inequality restrictions enhances the structural set identification, allowing for more precise inference, see Antolín-Díaz and Rubio-Ramírez, (2018). Moreover, in cases where impact sign restrictions alone cannot easily distinguish between different structural shocks, additional sign restrictions directly on the structural shocks provide a flexible way to separate them.
@@
Recently, the assumption that VAR errors follow a factor structure has gained popularity, with latent factors being interpreted as structural shocks (see, for instance, Korobilis, (2022), Banbura et al., (2023), Chan et al., (2022),Hauzenberger et al., (2022), Gambetti et al., (2023), Pfarrhofer and Stelzer, (2024), Prüser, (2024), Korobilis, (2025) and Chan and Qi, (2025)). We build on this literature, leveraging the factor structure for several key advantages. First, from a computational perspective, the factor structure enables equation-by-equation estimation, simplifying the estimation process and allowing for the inclusion of a rich set of variables in the model. Second, as the number of variables increases (including different measures of prices and output), it is reasonable to assume that the number of structural shocks is smaller than the number of variables, making a factor-based approach well-suited for high-dimensional settings. Third, and particularly important for our paper, the factors have a clear interpretation as structural shocks. This provides a natural framework for incorporating narrative restrictions in the form of prior distributions on the structural shocks, which allows to develop a simple yet efficient algorithm.
@@
Our approach deviates from existing methods such as those in Antolín-Díaz and Rubio-Ramírez, (2018) and Giacomini et al., (2021), which typically incorporate narrative restrictions through a pseudo-likelihood. Instead, imposing shock sign restrictions via prior distributions allows us to introduce a new sampling algorithm that offers significant computational efficiency by ensuring that each draw is accepted with certainty. This feature enables us to handle a large number of restrictions on both impact responses and structural shocks, overcoming computational challenges often encountered with accept-and-reject algorithms. Our approach therefore represents a major advancement in the scalability and feasibility of SVAR models, particularly in high-dimensional settings where existing algorithms may be very slow or even infeasible.
@@
In an empirical application, we demonstrate the practical utility of our approach by identifying five structural shocks in a large VAR. In this respect, we resemble Stock and Watson, (2012) and Hou, (2024) and estimate 30 variable VAR and disentangle structurally interpretable shocks by means of impact and shock sign restrictions. Importantly, we show that shock sign restrictions can help to separate structural shocks and to overall sharpen identification. In our empirical illustration, we provide evidence suggesting that financial risk shocks are a key driver of business cycle dynamics. This example underscores the flexibility and scalability of our methodology, which is capable of handling high-dimensional data and a large number of inequality restrictions on the impact responses as well as on the structural shocks.
@@
The remainder of this paper is organized as follows. Section 2 lays out and discusses the econometric framework. Section 3 applies the model to the study of Stock and Watson, (2012) to identify and investigate the effects of five structural shocks on the economy. Section 4 concludes.
@@ 2 A large structural VAR with factor structure
where $\bm{I}_{n}$ is the identity matrix of dimension $n$, $\otimes$ represents the Kronecker product, and $\bm{\beta}=\text{vec}([\bm{b}_{0},\bm{B}_{1},\dots,\bm{B}_{p}]^{\prime})$. The vector $\bm{x}_{t}=(1,\bm{y}^{\prime}_{t-1},\dots,\bm{y}^{\prime}_{t-p})^{\prime}$ of dimension $k=1+np$ contains an intercept and lagged values. The idiosyncratic component $\bm{v}_{t}$ accounts for measurement error or other idiosyncratic noise. In contrast the $r$ latent factors, $\bm{f}_{t}$ influencing multiple variables in the system and hence have the interpretation as structural shocks. More formally we multiply the model with the generalized inverse of $\bm{L}$ to obtain:
@@
Thus, $\bm{v}_{t}$ is treated as noise shocks without structural meaning, while $\bm{f}_{t}$ provides a projection of structural shocks in $\mathbb{R}^{r}$. Consequently, this formulation supports structural analysis using standard methods, including impulse response functions (e.g., Forni et al., (2019), Korobilis, (2022), Chan et al., (2022)).
Under the assumption of uncorrelated $\bm{f}_{t}$ and $\bm{v}_{t}$, the conditional variance of $\bm{u}_{t}$ is given by:
@@
To separately identify the common and idiosyncratic components, we adopt the condition $r\leq(n-1)/2$ from Anderson and Rubin, (1956). Under this restriction, for any observationally equivalent model $(\bm{L},\bm{\Sigma})$, it holds that $\bm{L}\bm{L}^{\prime}=\bm{L}^{*}\bm{L}^{*^{\prime}}$ and $\bm{\Sigma}=\bm{\Sigma}^{*}$.
However, without additional restrictions, $\bm{L}$ is not identified. Any orthogonal matrix $\bm{Q}\in\mathcal{O}(r)$, where $\mathcal{O}(r)=\{\bm{Q}\in\mathbb{R}^{r\times r}:\bm{Q}\bm{Q}^{\prime}=\bm{I}_{m}\}$, generates an equivalent model via the transformation $\tilde{\bm{L}}\tilde{\bm{f}}_{t}=\bm{L}\bm{Q}\bm{Q}^{\prime}\bm{f}_{t}$. In this paper we achive set identification by placing inequality restrictions on $\L$, $\bm{f}$ as well as on the product of both.
@@ 2.1 Linear inequality restrictions via prior distributions
We consider inequality restriction on the $\bm{L}$, $\bm{f}_{t}$, and on their product. [^2] In many cases we have a strong consensus in economic theory about the signs of impulse responses at impact but not at longer horizons (see, e.g., Canova and Paustian, (2011)). We can express these inequality restrictions as follows:
@@
where $\bm{l}_{i}$ denote the elements of $\bm{L}$ in the $i-$th equation, $\bm{R}_{i}^{L}$ is $r_{i}^{L}\times r$, $\bm{R}_{\tilde{t}}^{f}$ is $r_{\tilde{t}}^{f}\times r$, $\bm{R}_{i\tilde{t}}^{Lf}$ is $r_{i\tilde{t}}^{Lf}\times r$ and $\tilde{t}\in t=1,\dots,T$. We assume that $\ell_{i}^{L}<\upsilon_{i}^{L}$, $\ell_{\tilde{t}}^{f}<\upsilon_{\tilde{t}}^{f}$ and $\ell_{i\tilde{t}}^{Lf}<\upsilon_{i\tilde{t}}^{Lf}$. Furthermore, we allow elements in the lower and upper bound to be $\infty$- ∞ or $\infty$ to indicate that some inequity restrictions are single-sided. In many applications $\bm{R}_{i}^{L}$, $\bm{R}_{\tilde{t}}^{f}$ and $\bm{R}_{i\tilde{t}}^{Lf}$ will be equal to $\bm{I}_{r}$. The first set of inequality restrictions can be used to implement sign restrictions on $\bm{L}$ as in Korobilis, (2022). The second set of inequality restriction can be used to implement sign restrictions on the structural shocks $\bm{f}_{t}$ at specific time periods as in Antolín-Díaz and Rubio-Ramírez, (2018). Finally, the last set of inequality restrictions can be used to implement a magnitude restrictions on the product of both factors loadings and structural shocks as in Badinger and Schiman, 2023b . In particular, Badinger and Schiman, 2023b impose that one particular shock explains more then half of the unexplained movement of a certain variable at a specific period. This provides a flexible framework for the identification of structural shocks. Instead of only relaying on impact sign restrictions, researchers can also use shock sign restrictions to distinguish between different structural shocks. Or even use a single magnitude restriction as in Badinger and Schiman, 2023b to separate one shock from the others.
@@
where $\tilde{T}$ denotes the set of periods in which a restrictions is specified. We complete our model specification by describing the further prior distributions. For the variance terms of the noise we assume $\sigma_{j}^{2}\sim\mathcal{IG}(\alpha_{0},\beta_{0})$. In our empirical application, we set $\bm{l}_{0,i}=\bm{0}$, $\bm{V}_{\bm{l}_{i}}=10\times\bm{I}_{r}$ and $\alpha_{0}=\beta_{0}=0$. In high-dimensional settings such as large VARs, it is important to impose shrinkage prior to avoid overfitting. We follow Korobilis, (2022) and use the horseshoe prior proposed by Carvalho et al., (2010). Let $\bm{\beta}_{i}$ the VAR coefficients in the $i-$th equation, $i=1,\dots,n$ and $\beta_{i,j}$, the $j-$th coefficient in the $i-$th equation. Consider the prior for $\beta_{i,j}$, $i=1,\dots,n$ and $j=2,\dots,k$:
@@ 2.2 Gibbs Sampler
In this section, we develop an efficient posterior sampler to estimate the model. To impose the inequality restrictions on the factors or factor loadings we use the sampler from Botev, (2017). In contrast to existing rejecting sampling approaches (see, e.g. Antolín-Díaz and Rubio-Ramírez, (2018) or Giacomini et al., (2021) ) each draw is accepted with certainty. Posterior draws can be obtained by sampling sequentially from the conditional distributions:
@@
    $p(\bm{L}|\bm{y}\bm{\beta},,\bm{f},\bm{\Sigma},\bm{\lambda},\bm{\psi},\bm{z}_{\lambda},\bm{z}_{\bm{\psi}})=\prod_{i=1}^{n}=p(\bm{l}_{i}|\bm{y}_{i},\bm{\beta}_{i},\bm{f},\bm{\sigma}^{2}_{i})$
@@
$p(\psi_{i,j}|\beta_{i,j},\lambda_{i},z_{\psi_{i,j}})$ $\propto\psi_{i,j}^{\frac{1}{2}}\text{e}^{-\frac{1}{2\lambda_{i}\psi_{i,j}}(\beta_{i,j}-m_{i,j})^{2}}\times\psi^{-\frac{3}{2}}\text{e}^{-\frac{1}{\psi_{i,j}z_{\psi_{i,j}}}}$
$=\psi_{i,j}^{-2}\text{e}^{-\frac{1}{\psi_{i,j}}\left(\frac{1}{z_{\psi_{i,j}}}+\frac{(\beta_{i,j}-m_{i,j})^{2}}{2\lambda_{i}}\right)},$
@@
$(\psi_{i,j}|\beta_{i,j},\lambda_{i},z_{\psi_{i,j}})\sim\mathcal{IG}\left(1,\frac{1}{z_{\psi_{i,j}}}+\frac{(\beta_{i,j}-m_{i,j})^{2}}{2\lambda_{i}}\right).$ (23)
@@
$p(\lambda_{i}|\bm{\beta},\bm{\psi},z_{\lambda_{i}})$ $\propto\prod_{i}\lambda_{i}^{-\frac{1}{2}}\text{e}^{-\frac{1}{2\lambda_{i}\psi_{i,j}}(\beta_{i,j}-m_{i,j})^{2}}\times\lambda_{i}^{-\frac{3}{2}}\text{e}^{-\frac{1}{\lambda_{i}z_{\lambda_{i}}}},$
$=\lambda_{i}^{-\left(\frac{k}{2}+1\right)}\text{e}^{-\frac{1}{\lambda_{i}}\left(\frac{1}{z_{\lambda_{i}}}+\sum_{j=2}^{k}\frac{(\beta_{i,j}-m_{i,j})^{2}}{2\psi_{i,j}}\right)},$
@@
$(\lambda_{i}|\bm{\beta},\bm{\psi},z_{\psi_{\lambda_{i}}})\sim\mathcal{IG}\left(\frac{k}{2},\frac{1}{z_{\lambda_{i}}}+\sum_{j=2}^{k}\frac{(\beta_{i,j}-m_{i,j})^{2}}{2\psi_{i,j}}\right).$ (24)

### v2

@@ Abstract
We propose a high-dimensional structural vector autoregression framework that features a factor structure in the error terms and accommodates a large number of linear inequality restrictions on impact impulse responses, structural shocks, and their element-wise products. In particular, we demonstrate that narrative restrictions can be imposed via constraints on the structural shocks, which can be used to sharpen inference and disentangle structurally interpretable shocks. To estimate the model, we develop a highly efficient sampling algorithm that scales well with both the model dimension and the number of inequality restrictions on impact responses and structural shocks. It remains computationally feasible even in settings where existing algorithms may break down. To illustrate the practical utility of our approach, we identify five structural shocks and examine the dynamic responses of thirty macroeconomic variables, highlighting the model’s flexibility and feasibility in complex empirical applications. We provide empirical evidence that financial shocks are the most important driver of business cycle dynamics.
@@ 1 Introduction
Uhlig, (2005) popularizes the use of sign restrictions on impulse response functions, offering a flexible and theory-consistent approach to identifying structural shocks in Vector Autoregression (VAR) models. Instead of relying on often implausible zero restrictions, sign restrictions constrain the signs of impulse responses based on economic intuition. Building on this framework, Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018) introduce narrative restrictions, which incorporate external prior information about specific historical episodes-such as well-documented policy interventions or economic events-to further sharpen set identification. [^1] The sign restrictions on shocks over distinct time periods may be interpreted as a more agnostic framework than impact sign restrictions, as noted in Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018). Nonetheless, it is emphasized that sign restrictions must be grounded in a plausible narrative foundation. By conditioning structural shocks on particular time periods, narrative restrictions enable researchers to embed economically meaningful prior knowledge directly into the model, enhancing interpretability and credibility of the identified shocks. Using narrative restrictions have become popular recently. [^2] Important examples that use narrative restrictions include Zeev, (2018), Altavilla et al., (2019), Furlanetto and Robstad, (2019), Cheng and Yang, (2020), Kilian and Zhou, (2020), Kilian and Zhou, (2022), Laumer, (2020), Redl, (2020), Zhou, (2020), Antolin-Diaz et al., (2021), Caggiano et al., (2021), Larsen, (2021), Ludvigson et al., (2021), Maffei-Faccioli and Vella, (2021), Berger et al., (2022), Fanelli and Marsi, (2022), Inoue and Kilian, (2022), Badinger and Schiman, 2023a , Berthold, (2023), Caggiano and Castelnuovo, (2023), Conti et al., (2023), Harrison et al., (2023), Herwartz and Wang, (2023) Neri, (2023), Reichlin et al., (2023), Ascari et al., (2023), Boer et al., (2024) and Rüth and Van der Veken, (2023). However, in most applications researcher use rather small dimensional VARs with narrative restrictions. Furthermore, the implementation of narrative restriction rely on accept-and-reject algorithms (see, e.g., Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018) or Giacomini et al., (2021)), which may become computationally burdensome as the number of restrictions increases. Therefore, applications with narrative restriction are limited to VARs with a small number of narrative restrictions, typically for just a single structural shock.
@@
In this paper, we propose a high-dimensional Structural Vector Autoregression (SVAR) framework capable of accommodating a large number of linear inequality constraints on impact impulse responses, structural shocks, and their element-wise products. While previous literature has primarily focused on small narrative SVAR models, our approach extends these methods by enabling the inclusion of a much larger set of restrictions as well as a greater number of variables. The ability to incorporate a broader set of inequality restrictions enhances the structural set identification, allowing for more precise inference, see Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018). Moreover, in cases where impact sign restrictions alone cannot easily distinguish between different structural shocks, additional sign restrictions directly on the structural shocks provide a flexible way to separate them.
@@
Recently, the assumption that VAR errors follow a factor structure has gained popularity, with latent factors being interpreted as structural shocks (see, for instance, Korobilis, (2022), Chan et al., (2022), Hauzenberger et al., (2022), Banbura et al., (2023), Gambetti et al., (2023), Pfarrhofer and Stelzer, (2024), Prüser, (2024), Korobilis, (2025) and Chan and Qi, (2025)). We build on this literature, leveraging the factor structure for several key advantages. First, from a computational perspective, the factor structure enables equation-by-equation estimation, simplifying the estimation process and allowing for the inclusion of a rich set of variables in the model. Second, as the number of variables increases (including different measures of prices and output), it is reasonable to assume that the number of structural shocks is smaller than the number of variables, making a factor-based approach well-suited for high-dimensional settings. Third, and particularly important for our paper, the factors have a clear interpretation as structural shocks. This provides a natural framework for incorporating narrative restrictions in the form of prior distributions on the structural shocks, which allows to develop a simple yet efficient algorithm.
@@
Our approach deviates from existing methods such as those in Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018) and Giacomini et al., (2021), which typically incorporate narrative restrictions through a pseudo-likelihood. Instead, imposing shock sign restrictions via prior distributions allows us to introduce a new sampling algorithm that offers significant computational efficiency by ensuring that each draw is accepted with certainty. This feature enables us to handle a large number of restrictions on both impact responses and structural shocks, overcoming computational challenges often encountered with accept-and-reject algorithms. Our approach therefore represents a major advancement in the scalability and feasibility of SVAR models, particularly in high-dimensional settings where existing algorithms may be very slow or even infeasible. We compare our proposed algorithm with the approach of Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018) in a synthetic exercise in which we estimate several ten-variable VARs with different sets of impact and shock sign restrictions. The results indicate that our algorithm achieves substantially faster computation times than the method proposed by Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018) across the different identification schemes.
@@
In an empirical application, we demonstrate the practical utility of our approach by identifying five structural shocks in a large VAR. In this respect, we follow Stock and Watson, (2012) and estimate 30 variable VAR and disentangle structurally interpretable shocks by means of impact and shock sign restrictions. [^3] Specifically, we use the same set of 30 variables and identify the same structural shocks as in Stock and Watson, (2012). Hou, (2024) estimates the same 30-variables VAR and identifies the same 5 structural shocks. Importantly, we show that shock sign restrictions can help to separate structural shocks and to overall sharpen identification. In our empirical illustration, we provide evidence suggesting that financial risk shocks are a key driver of business cycle dynamics. This example underscores the flexibility and scalability of our methodology, which is capable of handling high-dimensional data and a large number of inequality restrictions on the impact responses as well as on the structural shocks.
@@
The remainder of this paper is organized as follows. Section 2 lays out and discusses the econometric framework. Section 3 studies the computational efficiency. Section 4 applies the model to the study of Stock and Watson, (2012) to identify and investigate the effects of five structural shocks on the economy. Section 4 concludes.
@@ 2 A large structural VAR with factor structure
where $\bm{I}_{n}$ is the identity matrix of dimension $n$, $\otimes$ represents the Kronecker product, and $\bm{\beta}=\text{vec}([\bm{b}_{0},\bm{B}_{1},\dots,\bm{B}_{p}]^{\prime})$. The vector $\bm{x}_{t}=(1,\bm{y}^{\prime}_{t-1},\dots,\bm{y}^{\prime}_{t-p})^{\prime}$ of dimension $k=1+np$ contains an intercept and lagged values. The idiosyncratic component $\bm{v}_{t}$ accounts for measurement error or other idiosyncratic noise. In contrast, the $r$ latent factors, $\bm{f}_{t}$ impact multiple variables in the system and hence have the interpretation as structural shocks. More formally we multiply the model with the generalized inverse of $\bm{L}$ to obtain:
@@
Thus, $\bm{v}_{t}$ is treated as noise shocks without structural meaning, while $\bm{f}_{t}$ provides a projection of structural shocks in $\mathbb{R}^{r}$. Consequently, this formulation supports structural analysis using standard methods, including impulse response functions (e.g., Forni et al., (2019), Korobilis, (2022), Chan et al., (2022)). Under the assumption of uncorrelated $\bm{f}_{t}$ and $\bm{v}_{t}$, the conditional variance of $\bm{u}_{t}$ is given by:
@@
To separately identify the common and idiosyncratic components, we adopt the condition $r\leq(n-1)/2$ from Anderson and Rubin, (1956). Under this restriction, for any observationally equivalent model $(\bm{L},\bm{\Sigma})$, it holds that $\bm{L}\bm{L}^{\prime}=\bm{L}^{*}\bm{L}^{*^{\prime}}$ and $\bm{\Sigma}=\bm{\Sigma}^{*}$. However, without additional restrictions, $\bm{L}$ is not identified. Any orthogonal matrix $\bm{Q}\in\mathcal{O}(r)$, where $\mathcal{O}(r)=\{\bm{Q}\in\mathbb{R}^{r\times r}:\bm{Q}\bm{Q}^{\prime}=\bm{I}_{m}\}$, generates an equivalent model via the transformation $\tilde{\bm{L}}\tilde{\bm{f}}_{t}=\bm{L}\bm{Q}\bm{Q}^{\prime}\bm{f}_{t}$. In this paper, we achieve set identification by placing inequality restrictions on $\bm{L}$ and $\bm{f}_{t}$. [^4] As shown below, our model also allows for inequality restrictions on the product of $\bm{L}$ and $\bm{f}$.
@@ 2.1 Linear inequality restrictions via prior distributions
We consider inequality restriction on the $\bm{L}$, $\bm{f}_{t}$, and on their product. [^5] In many cases we have a strong consensus in economic theory about the signs of impulse responses at impact but not at longer horizons (see, e.g. Canova and Paustian, (2011)). We can express these inequality restrictions as follows:
@@
where $\bm{l}_{i}$ denote the elements of $\bm{L}$ in the $i-$th equation, $\bm{R}_{i}^{L}$ is $r_{i}^{L}\times r$, $\bm{R}_{\tilde{t}}^{f}$ is $r_{\tilde{t}}^{f}\times r$, $\bm{R}_{i\tilde{t}}^{Lf}$ is $r_{i\tilde{t}}^{Lf}\times r$ and $\tilde{t}\in t=1,\dots,T$. We assume that $\ell_{i}^{L}<\upsilon_{i}^{L}$, $\ell_{\tilde{t}}^{f}<\upsilon_{\tilde{t}}^{f}$ and $\ell_{i\tilde{t}}^{Lf}<\upsilon_{i\tilde{t}}^{Lf}$. Furthermore, we allow elements in the lower and upper bound to be $\infty$- ∞ or $\infty$ to indicate that some inequity restrictions are single-sided. In many applications $\bm{R}_{i}^{L}$, $\bm{R}_{\tilde{t}}^{f}$ and $\bm{R}_{i\tilde{t}}^{Lf}$ will be equal to $\bm{I}_{r}$. The first set of inequality restrictions can be used to implement sign restrictions on $\bm{L}$ as in Korobilis, (2022). The second set of inequality restriction can be used to implement sign restrictions on the structural shocks $\bm{f}_{t}$ at specific time periods as in Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018). Finally, the last set of inequality restrictions can be used to implement a magnitude restrictions on the product of both factors loadings and structural shocks as in Badinger and Schiman, 2023b . In particular, Badinger and Schiman, 2023b impose that one particular shock explains more then half of the unexplained movement of a certain variable at a specific period. This provides a flexible framework for the identification of structural shocks. Instead of only relaying on impact sign restrictions, researchers can also use shock sign restrictions to distinguish between different structural shocks. Or even use a single magnitude restriction as in Badinger and Schiman, 2023b to separate one shock from the others. [^6] It should be noted that non-linear restrictions can be incorporated via an accept-rejection step. However, this may significantly increase the computational burden or even render sampling infeasible. If such non-linear restrictions can be well approximated by linear restrictions that fit within our framework, the accept-rejection step may become feasible and the additional computational cost can be kept minimal.
@@
where $\tilde{T}$ denotes the set of periods in which a restrictions is specified.
We complete our model specification by describing the further prior distributions. For the variance terms of the noise we assume $\sigma_{j}^{2}\sim\mathcal{IG}(\alpha_{0},\beta_{0})$. In our empirical application, we set $\bm{l}_{0,i}=\bm{0}$, $\bm{V}_{\bm{l}_{i}}=10\times\bm{I}_{r}$ and $\alpha_{0}=\beta_{0}=0$. In high-dimensional settings such as large VARs, it is important to impose shrinkage prior to mitigate the risk of overfitting. We follow Korobilis, (2022) and use the horseshoe prior proposed by Carvalho et al., (2010). Let $\bm{\beta}_{i}$ the VAR coefficients in the $i-$th equation, $i=1,\dots,n$ and $\beta_{i,j}$, the $j-$th coefficient in the $i-$th equation. Consider the prior for $\beta_{i,j}$, $i=1,\dots,n$ and $j=2,\dots,k$:
@@ 2.2 Gibbs Sampler
In this section, we develop an efficient posterior sampler to estimate the model. To impose the inequality restrictions on the factors or factor loadings we use the sampler from Botev, (2017). In contrast to existing rejecting sampling approaches (see, e.g. Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018) or Giacomini et al., (2021)) each draw is accepted with certainty. Posterior draws can be obtained by sampling sequentially from the conditional distributions:
@@
    $p(\bm{L}|\bm{y}\bm{\beta},\bm{f},\bm{\Sigma},\bm{\lambda},\bm{\psi},\bm{z}_{\lambda},\bm{z}_{\bm{\psi}})=\prod_{i=1}^{n}=p(\bm{l}_{i}|\bm{y}_{i},\bm{\beta}_{i},\bm{f},\bm{\sigma}^{2}_{i})$
@@
$p(\psi_{i,j}|\beta_{i,j},\lambda_{i},z_{\psi_{i,j}})$ $\propto\psi_{i,j}^{\frac{1}{2}}\text{e}^{-\frac{1}{2\lambda_{i}\psi_{i,j}}(\beta_{i,j})^{2}}\times\psi^{-\frac{3}{2}}\text{e}^{-\frac{1}{\psi_{i,j}z_{\psi_{i,j}}}}$
$=\psi_{i,j}^{-2}\text{e}^{-\frac{1}{\psi_{i,j}}\left(\frac{1}{z_{\psi_{i,j}}}+\frac{(\beta_{i,j})^{2}}{2\lambda_{i}}\right)},$
@@
$(\psi_{i,j}|\beta_{i,j},\lambda_{i},z_{\psi_{i,j}})\sim\mathcal{IG}\left(1,\frac{1}{z_{\psi_{i,j}}}+\frac{(\beta_{i,j})^{2}}{2\lambda_{i}}\right).$ (23)
@@
$p(\lambda_{i}|\bm{\beta},\bm{\psi},z_{\lambda_{i}})$ $\propto\prod_{i}\lambda_{i}^{-\frac{1}{2}}\text{e}^{-\frac{1}{2\lambda_{i}\psi_{i,j}}(\beta_{i,j})^{2}}\times\lambda_{i}^{-\frac{3}{2}}\text{e}^{-\frac{1}{\lambda_{i}z_{\lambda_{i}}}},$
$=\lambda_{i}^{-\left(\frac{k}{2}+1\right)}\text{e}^{-\frac{1}{\lambda_{i}}\left(\frac{1}{z_{\lambda_{i}}}+\sum_{j=2}^{k}\frac{(\beta_{i,j})^{2}}{2\psi_{i,j}}\right)},$
@@
$(\lambda_{i}|\bm{\beta},\bm{\psi},z_{\psi_{\lambda_{i}}})\sim\mathcal{IG}\left(\frac{k}{2},\frac{1}{z_{\lambda_{i}}}+\sum_{j=2}^{k}\frac{(\beta_{i,j})^{2}}{2\psi_{i,j}}\right).$ (24)
@@
3 Computational Efficiency
This section evaluates the computational performance of our proposed algorithm in comparison to the approach developed by Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018) (henceforth DDR18). To assess the relative efficiency, we consider three different structural VAR specifications with $n=10$ variables, lag length $p=4$, $k=5$ structural shocks, and a sample size of $T=148$ observations. [^7] In our empirical application in Section 4, we use quarterly data from 1983Q1 to 2019Q4, corresponding to $T=148$. We identify $k=5$ structural shocks and set the lag length to $p=4$, consistent with the simulation design here.
[…]
Table 1: Comparison of DDR18 algorithm and proposed algorithm
Restrictions # Restrictions Code Mean
Impact Sign 15 DRR18 32.09
Shock Sign 0 Proposed 0.18
Impact Sign 15 DRR18 33.32
Shock Sign 3 Proposed 0.23
Impact Sign 15 DRR18 34.58
Shock Sign 6 Proposed 0.23
Notes: The table reports the average runtime (in minutes) required by the DDR18 algorithm and our proposed algorithm to obtain 100 admissible draws across three different structural VAR specifications. The # Restriction column indicates the number of imposed impact and shock sign restrictions.
Table 1 summarizes the computational comparison between the proposed algorithm and DDR18 across the three VAR configurations. It is evident that the proposed algorithm achieves substantially faster runtimes in all cases. As highlighted in Footnote 8, the DDR18 algorithm was originally developed for small-scale VARs, whereas our approach is explicitly designed for high-dimensional VARs with factor error structures, which requires more sophisticated shrinkage methods. Importantly, in our simulation design, we construct the data-generating process such that the imposed shock sign restrictions align with the true signs of the underlying structural shocks. In the absence of this feature, the DDR18 algorithm fails to find a sufficient number of admissible draws. [^9] See Appendix Data Generation Process for a detailed discussion of the data-generating process. Our computations were carried out using MATLAB R2024b on an Intel Core Ultra 7 with a 1.70 GHz base speed and 8 GB of RAM.
Table 1 clearly shows that the proposed algorithm achieves substantially faster runtimes than the DDR18 algorithm. Notably, the runtime of the proposed method remains stable even as the number of shock sign restrictions increases. In contrast, as the number of variables $n$ or the number of identifying restrictions grows - e.g., to $n=20$ or $n=30$ with more impact and shock sign restrictions - the DDR18 algorithm struggles or even fails to find admissible draws within a feasible timeframe. These results underscore the computational advantages of our approach and highlight its suitability for estimating large-scale VARs with extensive identifying restrictions. In the empirical application that follows, we apply the proposed algorithm to a thirty-variable VAR identified with 60 impact sign restrictions and 26 shock sign restrictions.